# An Optimal Design Framework for Lasso Sign Recovery

### v2

@@ 5.2 Scenario 2: $n=9$, $p=10$ Known Signs
Figure 5 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2022), and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings. To demonstrate the added value these three designs have by incorporating sign information, the figure also has the $\Phi_{\lambda}$ curve for a $UE(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. For all $\beta$ and $k$ settings except $\beta=4$ and $k=5$, the HILS-optimal design shows a larger $\Phi_{\text{max}}$ and area under the curve compared to DCD and $Var(s+)$. For $\beta=4$ and $k=5$, Figure 5 shows a slightly larger $\Phi_{\text{max}}$ for DCD compared to the HILS-optimal design. However, the HILS-optimal design sees higher $\Phi_{\lambda}$ for larger values of $\log(\lambda)$ (around $\log(\lambda)=1$) and has a larger area under its $\Phi_{\lambda}$ curve. Thus, we deem that the HILS-optimal design is superior to the competitors in all settings demonstrated in Figure 5. Finally, we note the poor performance of the $UE(s^{2})$-optimal design compared to all designs, especially when $k=5$ with $\Phi_{\lambda}\approx 0$.
Figure 5: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 2.
@@ 5.3 Scenario 3: $n=14$, $p=20$ Known Signs
Figure 6 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2022), and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings. We again included the $\Phi_{\lambda}$ curve for a $UE(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. Across all $\beta$ and $k$ settings there is very little difference between the performances of the three designs that include sign information. However, all of these designs had significantly higher $\Phi_{\lambda}$ values than the $UE(s^{2})$-optimal design.
Figure 6: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 3.
The lack of separation between the HILS-optimal design from the other designs is not unexpected. The larger the discrepancy between $p$ and $n$, the less we expect $||\textbf{C}-\textbf{C}^{*}||_{F}$ to be a good measure because we are dealing with a scenario in which C will never be close to compound symmetric. In scenarios 1 and 2, $n$ is close to $p$, and, thus, C can be close to the ideal structure of compound symmetry. Nonetheless, the HILS-optimal design was found much more efficiently than the DCD which took on the order of dozens of hours. Since the HILS-optimal design can be constructed more quickly and show comparable performance to the DCDs, we view this algorithmic design construction approach as worthwhile.
@@ 6 Summary
The SSD literature has predominately constructed designs by optimizing heuristic criteria that measure a design’s proximity to a (nonexistent) orthogonal design. The heuristic criteria are tractable in their optimization, in that optimal designs can generally be constructed directly or algorithmically in a reasonable amount of time. However, these heuristic criteria are not directly tied to a screening analysis method, so there is no guarantee that the resulting analysis under an optimal design will have good statistical properties. This article resolves this disconnect by optimizing criteria based on the probability of sign recovery under the lasso, which is well-defined even when $n-1<p$. Our major contributions are:
@@
    A local optimality criterion assuming known $\boldsymbol{\beta}$ and fixed $\lambda$. A trivial design that confounds all inactive factors with the intercept is shown to be optimal for sign recovery. An exact design construction is given that can improve the probability of sign recovery over an orthogonal design for some $\lambda$ The design has positive and nearly constant pairwise correlations, following the ideal structure for the $Var(s+)$-criterion.
@@
    A rigorous proof that across all completely symmetric matrices, an orthogonal correlation matrix is suboptimal for nearly all $\lambda>0$ when $\boldsymbol{z}$ is known. The optimal form instead takes on positive, constant off-diagonal elements. In the case of unknown sign information, an orthogonal correlation matrix is shown to be a local maximum for a range of $\lambda$. These optimal forms rigorously justify the $UE(s^{2})$- and $Var(s+)$-criterion in the cases of unknown and known $\boldsymbol{z}$, respectively. This at least partially solves the two open problems from Section 1.1.
@@
Our work provides an alternative to using simulation to compare SSDs, which can be tedious and difficult to reproduce independently. In particular, there are at least two reasons to be skeptical of conclusions drawn from simulations in the context of SSDs. First, simulation can be misleading when subtly different versions of complicated statistical procedures are used. In our attempts to reproduce and compare the simulation studies of Singh and Stufken (2022) and Weese et al. (2021), we discovered the regularization methods used were sensitive to two different aspects of the procedure implementation. In particular, these papers used the Gauss-Dantzig selector with different ways of exploring the tuning parameter space. They also differed in how they implemented a secondary thresholding of the estimates which determines a subset of the active factors that should remain active. Typically this threshold removes estimates whose absolute magnitude are less than $\sigma^{2}$, but thresholding in this way is of dubious reliability because there is no natural way to estimate $\sigma^{2}$ and thus threshold levels are more or less arbitrary. The approach we present in this paper avoids these difficulties by eliminating the need for simulation at all, in a similar way that closed-form power analysis procedures are routinely used in simple, replicated experimental design settings.
@@
Another danger of the existing heuristic criteria is that there is no guarantee the optimal design has any statistical value. For example, there technically exists a $UE(s^{2})$-optimal design with $n=2$ runs and $p=100$ factors, but this design has no statistical value. Our criteria, however, do reflect statistical value and so can be used to give objective information about design quality, akin to a power analysis. Experimenters can specify a minimum effect size of interest, $\beta$, as well as an educated guess regarding the number of factors that are likely to have such an effect size and investigate the sign-recovery probabilities as a function of $\lambda$. If the maximum average probability is close to 0, it is unlikely that the design can provide reliable sign recovery for the specified effect size. Such a low value could lead to a reconsideration of the runs budget, expected sparsity, and/or the size of effects that are deemed of interest to detect.
@@
Our framework allows an even deeper investigation, if desired. For example, we can determine whether a low average probability is due to an inability to reliably exclude all inactive effects (the $I_{\lambda}$ event) or to reliably include all active effects (the $S_{\lambda}$ event). In screening experiments, we are much more concerned with identifying the true effects than with allowing inactive effects through the filter. This suggests that the $S_{\lambda}$ event is more important to the experimenter than the $I_{\lambda}$ event.
There are many avenues of future research to be explored. First, it is important to establish the concavity (or log concavity) of the criteria that are summarized over $\lambda$. We have already seen that such criteria are not concave for a fixed $\lambda$. Second, new construction techniques for designs approaching the desired structure of the $Var(s+)$-criterion would be incredibly useful. Continuing with exact design construction, while our algorithm is faster relative to Singh and Stufken (2022), it is much slower than that for the heuristic criteria. More work is needed to make the design construction algorithms faster. Finally, we are currently extending the criteria to include the case where inactive factors have small, but nonzero, effects as well as considering the case of criteria for a thresholded lasso.
@@
Following the References are proofs and computational details on the evaluation of the criteria. R files that construct designs following (14); identify optimal $c$ values for the criteria in Section 3; and construct designs with the HILS algorithm. may be found at https://github.com/hkyoung361/Lasso_Optimal_SSD. Code is also included for the examples in Section 5.

### v3

@@ 5.2 Scenario 2: $n=9$, $p=10$ Known Signs
Figure 8 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the CS-optimal design, HILS-optimal design, DCD (Singh and Stufken, 2023), and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings. To demonstrate the added value these three designs have by incorporating sign information, the figure also has the $\Phi_{\lambda}$ curve for a $\text{UE}(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. For all $\beta$ and $k$ settings except $\beta=4$ and $k=5$, the HILS-optimal design shows a larger $\Phi_{\text{max}}$ and area under the curve compared to DCD and $\text{Var}(s+)$. For $\beta=4$ and $k=5$, Figure 8 shows a slightly larger $\Phi_{\text{max}}$ for DCD compared to the HILS-optimal design. However, the HILS-optimal design sees higher $\Phi_{\lambda}$ for larger values of $\log(\lambda)$ (around $\log(\lambda)=1$) and has a larger area under its $\Phi_{\lambda}$ curve. Thus, we deem that the HILS-optimal design is superior to the competitors in all settings demonstrated in Figure 8. Finally, we note the poor performance of the $\text{UE}(s^{2})$-optimal design compared to all designs, especially when $k=5$ with $\Phi_{\lambda}\approx 0$. This is most likely due to the prevalence of negative correlations leading to singular $C_{\mathcal{A}}$ for some supports.
Figure 8: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 2. The HILS optimal design shown in all panels was chosen assuming $k=3$ and $\beta=3$. The CS Optimal curve refers to the hypothetical design having the optimal completely symmetric form for $k=3$ and $\beta=3$.
@@ 5.3 Scenario 3: $n=14$, $p=20$ Known Signs
Figure 9 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2023), and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings. We again included the $\Phi_{\lambda}$ curve for a $\text{UE}(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. Across all $\beta$ and $k$ settings there is very little difference between the performances of the three designs that include sign information. However, all of these designs had significantly higher $\Phi_{\lambda}$ values than the $\text{UE}(s^{2})$-optimal design.
Figure 9: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 3. The HILS optimal design shown in all panels was chosen assuming $k=5$ and $\beta=2$. The CS Optimal curve refers to the hypothetical design having the optimal completely symmetric form for $k=5$ and $\beta=2$.
The lack of separation between the HILS-optimal design from the other designs was not unexpected given $n$ is nearly half of $p$. It is challenging to find a design that simultaneously minimizes correlations. Nonetheless, the HILS-optimal design performed comparably to the DCD and $\text{Var}(s+)$-optimal designs, with all three significantly outperforming the $\text{UE}(s^{2})$-optimal design. The HILS-optimal design was found much faster than the DCD, which took us dozens of hours to construct. Since the HILS-optimal design can be constructed more quickly and show comparable performance to the DCDs, we view this algorithmic design construction approach as worthwhile.
@@ 6 Summary
The SSD literature has predominately constructed designs by optimizing heuristic criteria that measure a design’s proximity to a (nonexistent) orthogonal design. The heuristic criteria are tractable in their optimization, in that optimal designs can generally be constructed directly or algorithmically in a reasonable amount of time. However, these heuristic criteria are not directly tied to a screening analysis method, so there is no guarantee that the resulting analysis under an optimal design will have good statistical properties. This article resolves this disconnect by optimizing criteria based on the probability of sign recovery under the lasso, which is well-defined even when $n-1<p$. The major contributions are:
@@
    A local optimality criterion assuming known $\boldsymbol{\beta}$ and fixed $\lambda$. A trivial design that confounds all inactive factors with the intercept is shown to be optimal for sign recovery. An exact design construction is given that can improve the probability of sign recovery over an orthogonal design for some $\lambda$ The design has positive and nearly constant pairwise correlations, following the ideal structure for the $\text{Var}(s+)$-criterion.
@@
    A rigorous proof that across all completely symmetric matrices, an orthogonal correlation matrix is suboptimal for nearly all $\lambda>0$ when $\boldsymbol{z}$ is known. The optimal form instead takes on positive, constant off-diagonal elements. In the case of unknown sign information, an orthogonal correlation matrix is shown to be a local maximum for a range of $\lambda$. These optimal forms rigorously justify the $\text{UE}(s^{2})$- and $\text{Var}(s+)$-criterion in the cases of unknown and known $\boldsymbol{z}$, respectively. This at least partially solves the two open problems from Section 1.1.
@@
The proposed optimal design framework provides an alternative to using simulation to compare SSDs, which can be tedious and difficult to reproduce independently. In particular, there are at least two reasons to be skeptical of conclusions drawn from simulations in the context of SSDs. First, simulations can be misleading when subtly different versions of complicated statistical procedures are used. In our attempts to reproduce and compare the simulation studies of Singh and Stufken (2023) and Weese et al. (2021), we discovered the regularization methods used were sensitive to two different aspects of the procedure implementation. In particular, these papers used the Gauss-Dantzig selector with different ways of exploring the tuning parameter space. They also differed in how they implemented a secondary thresholding of the estimates which determines a subset of the active factors that should remain active. Typically this threshold removes estimates whose absolute magnitude are less than $\sigma^{2}$, but thresholding in this way is of dubious reliability because there is no natural way to estimate $\sigma^{2}$ and thus threshold levels are more or less arbitrary. The approach we present in this paper avoids these difficulties by eliminating the need for simulation at all, in a similar way that closed-form power analysis procedures are routinely used in simple, replicated experimental design settings.
@@
Another danger of the existing heuristic criteria is that there is no guarantee the optimal design has any statistical value. For example, there technically exists a $\text{UE}(s^{2})$-optimal design with $n=2$ runs and $p=100$ factors, but this design has no statistical value. The lasso criteria, however, do reflect statistical value and so can be used to give objective information about design quality, akin to a power analysis. Experimenters can specify a minimum effect size of interest, $\beta$, as well as an educated guess regarding the number of factors that are likely to have such an effect size and investigate the sign-recovery probabilities as a function of $\lambda$. If the maximum average probability is close to 0, it is unlikely that the design can provide reliable sign recovery for the specified effect size. Such a low value could lead to a reconsideration of the runs budget, expected sparsity, and/or the size of effects that are deemed of interest to detect.
@@
The proposed framework allows an even deeper investigation, if desired. For example, we can determine whether a low average probability is due to an inability to reliably exclude all inactive effects (the $I_{\lambda}$ event) or to reliably include all active effects (the $S_{\lambda}$ event). In screening experiments, we are much more concerned with identifying the true effects than with allowing inactive effects through the filter. This suggests that the $S_{\lambda}$ event is more important to the experimenter than the $I_{\lambda}$ event.
There are many avenues of future research to be explored. First, we noted in multiple sections and examples that for larger $\lambda$, $\Phi^{\pm}_{\lambda}(\textbf{C}\,|\,\boldsymbol{\beta})$ sometimes is optimized by a C with large positive off-diagonals, but its improvement is due solely to its ability to recover $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$. This is practically undesirable because we would like to do well across many possible sign vectors. A similar criticism also holds for averaging across all $\mathcal{A}\in\mathcal{A}_{k}$. That is, the best design may do great for only some supports, and poorly for others. It would be interesting to consider such criteria that consider the minimum probability rather than the average. However, we expect that such criteria would have low probabilities that they would not seem practically useful, as it focuses on the worst possible scenario.
Another area of future work is to explore direct construction techniques for designs with all positive correlations, as such designs may be strong candidates to optimize the known-sign criteria. Continuing with exact design construction, while the HILS algorithm is faster relative to Singh and Stufken (2023), it is much slower than that for the heuristic criteria. More work is needed to make faster design construction algorithms.
Finally, it is important to extend the criteria to incorporate more practical assumptions. Rather than assuming a known magnitude of the active effects, it would be better to incorporate uncertainty of the magnitudes using a prior distribution. This may require even heavier computation, as one would have to integrate the proposed criteria over the prior distribution. We are also considering criteria where the inactive factors have small, but nonzero, effects. This requires a definition of sign recovery with respect to the thresholded lasso in which estimated effects below some threshold value are set to 0.
@@
Following the References are proofs of all results as well as computational details on the evaluation of our criteria. R files that construct designs following (14); replicate the numerical studies in Section 3.1; identify optimal $c$ values for the criteria in Section 3; and construct designs with the HILS algorithm may be found at https://github.com/hkyoung361/Lasso_Optimal_SSD.